OpenStudy (anonymous):

find the integreal of (x-1)/(x^(1/2)) from 1 to 9. please help i have a math test coming up medals will be given

OpenStudy (istim):

Split the integral. x/ sqrt(x) - 1/sqrt(x) dx simplify. sqrt(x) - x^(-1/2) dx simple integration from there.

OpenStudy (anonymous):

how did it become sqrt of x

OpenStudy (istim):

Sorry, not "split the integral", split the equation. Splitting the integral is a different thing.

OpenStudy (istim):

sqrt(x) = x^(1/2)

OpenStudy (anonymous):

so we have x/sqrt(x)- 1/sqrtx

OpenStudy (anonymous):

and then how did you simplfy

OpenStudy (anonymous):

cross multiply?

OpenStudy (anonymous):

how is x/sqrt of x= sqrt of x, is what im really asking

OpenStudy (istim):

Exponent rules. x^(a)/ (x^(b)) = x^(a-b) http://2.bp.blogspot.com/_ViOqoLbfiLQ/THWS177DVwI/AAAAAAAAANE/1QP4DymkDi0/s1600/exponent+rules.png

OpenStudy (istim):

so x^(1) / (x^(1/2)) = x^ ( 1 - 1/2) = x^(1/2) which is just.. sqrt(x)

OpenStudy (anonymous):

oh my, i get it now! thank you so mcuh

OpenStudy (anonymous):

Istime can i ask you another question?